OpenStudy (anonymous):

In a rowboat, it took Alice 6 hours to go 6 miles upstream, but she was able to return to her starting point in only 45 minutes. Assuming her rowing speed is constant, what was the speed of the current?

OpenStudy (anonymous):

When you row at a constant rate, and the speed of the current is x, when you are going with the current your boat is actually moving at (constant rate + current) and when you are going against the current, your boat is actually going at (constant rate - current) d = rt 6 = (c - x)6 6 = (c + x)(.75) 45 minutes is 3/4 of an hour 6 = 6c - 6x distribute the 1st equation 6 = .75c + .75x distribute the 2nd equation Solve the 1st equation for c and then we are going to substitute. 6 + 6x = 6c divide thru by 6 1 + x = c substitute into the other equation 6 = .75(1 + x) + .75x 6 = .75 + .75x + .75x 6 = .75 + 1.5x 5.25 = 1.5x 3.5 = x Current is 3.5 mph

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

thankss